OpenStudy (anonymous):

A conveyor belt moves 450 ft in 3/8 h. Complete the unit rate. ft/h

OpenStudy (anonymous):

@KamiBug

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (blurbendy):

450 / 3/8 = 450 * 8/3 450 * 8/3 = 150 * 8 = ??? ft/h

OpenStudy (blurbendy):

understand?

OpenStudy (blurbendy):

@cjackson2002